# On the Cuspy Structure of Rotating Wormhole Shadows

### v1

@@ 5 Conclusion and discussion
By performing a two-parameter scan over $\lambda$ and $a$, we have constructed a detailed phase diagram (Fig. 5) that organizes the distinct shadow morphologies into four well-separated phases. Crucially, the formation of cusps, marking the transition between smooth and cuspy shadow boundaries, only becomes possible when $\lambda$ is allowed to vary. We have identified a critical value $\lambda_{c}=\frac{1}{4}(\sqrt{5}-1)$ at which the morphology of the shadow undergoes a qualitative transition. Remarkably, this critical value is independent of both the spin parameter $a$ and the throat radius $r_{0}$. For $\lambda<\lambda_{c}$, the outer circular orbit curve and the throat orbit curve meet smoothly with identical slope. No cusp or swallowtail appears. For $\lambda>\lambda_{c}$, the shadow exhibits a prominent cusp with slope discontinuity at the intersection point, clearly visible in Fig. 3. The outer unstable orbit curve develops a self-intersecting swallowtail structure. Working with fixed $\lambda>\lambda_{c}$, as the spin parameter is decreased, the two “ears” of the swallowtail merge, forming a closed loop; both orbit families still contribute to the boundary. For very small $a$, the throat orbit detaches completely from the shadow boundary, rendering the shadow solely determined by the outer unstable circular orbits. Within this phase, for sufficiently small spin (e.g., $a\leq 0.033$ at $\lambda=2$), the throat orbit further shrinks to a single point and ultimately vanishes, implying that no light signal from the throat can reach the distant observer. The ears-touching boundary exhibits a linear dependence between $a$ and $\lambda$, as shown in Fig. 5. The throat-drowning happens only when $a<0.08$. For higher-spinning wormholes, the throat invariably participates in the shadow.
In conclusion, we have demonstrated that the shadow of rotating traversable wormholes exhibits a surprisingly rich structure governed by the redshift parameter $\lambda$ and spin parameter $a$. The emergence of a cusp, the phase diagram comprising four distinct morphologies, and the re-entrant behavior of the throat orbit are novel features characterizing the wormhole shadow. This systematic morphological analysis deepens our theoretical understanding of wormhole spacetimes and offers a clean observational diagnostic for constraining $\lambda$ and $a$ with future high-resolution images.
The recent discovery [23] shows that cusp formation in black hole shadows constitutes a topological phase transition, characterized by a flip of the topological charge. Moreover, the cusp exhibits the equal-area law and universal critical exponents. The swallowtail shape of the outer orbit curve and the appearance/disappearance of the cusp for wormhole shadow strongly suggest a similar pattern. It is compelling to further study whether the same mechanism from [23] applies to the wormhole shadow or not.
@@ Acknowledgements
Appendix A Parameterization of the wormhole throat orbit
@@
The parametric equations for the throat orbit can be expressed as
@@
with $\bar{r}\in[0,1]$. $g(\bar{r})$ is the core of the parameterization. This parameterization should ensure the following properties, which put constraints on $g(\bar{r})$. When $\bar{r}=0$, it corresponds to the right endpoint of the ellipse $(\alpha_{c}+\mathcal{A},0)$. This condition demands $g(0)=1$. When $\bar{r}=1$, it corresponds to the impact parameters of the outer unstable orbit at $r_{\text{ph}}\to r_{0}$, denoted as
@@
Moreover, we usually ask the derivatives of $\alpha$ and $\beta$ with respect to $\hat{r}$, at $\hat{r}=1$, equals derivatives obtained from the outer unstable orbit at $r_{\text{ph}}\to r_{0}$. This condition further puts constraints on $g^{\prime}(\bar{r})|_{\bar{r}=1}$. Once the boundary condition for $\alpha(\bar{r})$ at $\bar{r}=1$ is satisfied, the boundary condition for $\beta(\bar{r})$ would be satisfied automatically, due to the compatibility condition. One can find a favorite parameterization for the wormhole throat in terms of $g(\bar{r})$, as far as the parameterization is consistent with the boundary conditions for $(\alpha(\bar{r}),\beta(\bar{r}))$.

### v2

@@ 5 Conclusion and discussion
By performing a two-parameter scan over $\lambda$ and $a$, we have constructed a detailed phase diagram (Fig. 5) that organizes the distinct shadow morphologies into four well-separated phases. Crucially, the formation of cusps, marking the transition between smooth and cuspy shadow boundaries, only becomes possible when $\lambda$ is allowed to vary. We have identified a critical value $\lambda_{c}=\frac{1}{4}(\sqrt{5}-1)$ at which the morphology of the shadow undergoes a qualitative transition. Remarkably, this critical value is independent of both the spin parameter $a$ and the throat radius $r_{0}$. For $\lambda<\lambda_{c}$, the outer circular orbits curve and the throat orbits curve meet smoothly with identical slope. No cusp or swallowtail appears. For $\lambda>\lambda_{c}$, the shadow exhibits a prominent cusp with slope discontinuity at the intersection point, clearly visible in Fig. 3. The outer unstable orbits curve develops a self-intersecting swallowtail structure. Working with fixed $\lambda>\lambda_{c}$, as the spin parameter is decreased, the two “ears” of the swallowtail merge, forming a closed loop; both orbit families still contribute to the boundary. For very small $a$, the throat orbits detach completely from the shadow boundary, rendering the shadow solely determined by the outer unstable circular orbits. Within this phase, for sufficiently small spin (e.g., $a\leq 0.033$ at $\lambda=2$), the throat orbits further shrink to a single point and ultimately vanish, implying that no light signal from the throat can reach the distant observer. The ears-touching boundary exhibits a linear dependence between $a$ and $\lambda$, as shown in Fig. 5. The throat-drowning happens only when $a<0.08$. For higher-spinning wormholes, the throat invariably participates in the shadow.
In conclusion, we have demonstrated that the shadow of rotating traversable wormholes exhibits a surprisingly rich structure governed by the redshift parameter $\lambda$ and spin parameter $a$. The emergence of a cusp, the phase diagram comprising four distinct morphologies, and the re-entrant behavior of the throat orbits are novel features characterizing the wormhole shadow. This systematic morphological analysis deepens our theoretical understanding of wormhole spacetimes and offers a clean observational diagnostic for constraining $\lambda$ and $a$ with future high-resolution images.
The recent discovery [39] shows that cusp formation in black hole shadows constitutes a topological phase transition, characterized by a flip of the topological charge. Moreover, the cusp exhibits the equal-area law and universal critical exponents. The swallowtail shape of the outer orbits curve and the appearance/disappearance of the cusp for wormhole shadow strongly suggest a similar pattern. It is compelling to further study whether the same mechanism from [39] applies to the wormhole shadow or not.
@@ Acknowledgements
Appendix A Parameterization of the wormhole throat orbits
@@
The parametric equations for the throat orbits can be expressed as
@@
with $\bar{r}\in[0,1]$. $g(\bar{r})$ is the core of the parameterization. This parameterization should ensure the following properties, which put constraints on $g(\bar{r})$. When $\bar{r}=0$, it corresponds to the right endpoint of the ellipse $(\alpha_{c}+\mathcal{A},0)$. This condition demands $g(0)=1$. When $\bar{r}=1$, it corresponds to the impact parameters of the outer unstable orbits at $r_{\text{ph}}\to r_{0}$, denoted as
@@
Moreover, we usually ask the derivatives of $\alpha$ and $\beta$ with respect to $\hat{r}$, at $\hat{r}=1$, equals derivatives obtained from the outer unstable orbits at $r_{\text{ph}}\to r_{0}$. This condition further puts constraints on $g^{\prime}(\bar{r})|_{\bar{r}=1}$. Once the boundary condition for $\alpha(\bar{r})$ at $\bar{r}=1$ is satisfied, the boundary condition for $\beta(\bar{r})$ would be satisfied automatically, due to the compatibility condition. One can find a favorite parameterization for the wormhole throat in terms of $g(\bar{r})$, as far as the parameterization is consistent with the boundary conditions for $(\alpha(\bar{r}),\beta(\bar{r}))$.